# Learning Interpretable Hierarchical Dynamical Systems Models from Time Series Data

### v1

@@ A.1.6 Hyperparameters
In Tab. 3 we report the hyperparameters used for training the shPLRNN models on the datasets in Sect. 4.1. These include the latent dimension $M$ (equivalent to the dimension of the observed data), the hidden dimension $L$, the GTF interpolation strength $\alpha$, as well as the number of features $N_{\text{feat}}$. As recommended in Hess et al. [2023], $\alpha$ decays exponentially from $\alpha_{\text{start}}$ at the beginning to $\alpha_{\text{end}}$ at the end of training.
@@ A.2 Evaluation Measures
In lower dimensional state spaces this may be approximated by discrete binning: Let $m$ be the number of bins in each dimension and $\hat{p}^{(k)}$ the number of points in the $k$-th bin, then with a total number of bins $N=m^{d}$, an estimate of the state space divergence is given by
$D_{\text{stsp}}\simeq\sum_{k=1}^{N}\hat{p}_{\text{gen}}^{(k)}(x)\log\frac{\hat{p}_{\text{gen}}^{(k)}(x)}{\hat{p}_{\text{true}}^{(k)}(x)}$ (10)
Since the total number of bins scales exponentially with the observation dimension $N$, for experiments with the Lorenz-96 system, we used $m=5$ per dimension, as suggested in Hemmer et al. [2024], to ensure the approach remained computationally feasible. For very high-dimensional spaces, Brenner et al. [2022], Mikhaeil et al. [2022] suggest a GMM-based estimate with Gaussians centered on the trajectory points (which was not required here, however). For the values in Table 1, we sampled single time series with $10,000$ time steps from both the ground truth system (serving as the test set) and from each trained model, inferring only the initial state from the first time step of the test set, and cutting off transients ($1,000$ time steps) to focus on the long term attractor behavior.
@@ A.4 Comparison Methods
As in LEADS, in CoDA (Context-Informed Dynamics Adaptation, Kirchmeyer et al. [2022]) the dynamics for each environment $e\in E$ is modeled as a combination of shared dynamics $f$ and environment-specific dynamics. However, here the environment-specific dynamics are modeled via a low-dimensional context vector $\xi_{e}$ that is used to generate environment-specific dynamics through a hypernetwork:
$f_{e}(x)=f(x)+W\xi_{e}$ (18)
where $W$ is a learned weight matrix. As for LEADS, the learning objective is to minimize the contribution of the context vectors $\xi_{e}$ while ensuring accurate modeling of the individual system dynamics. This is achieved by solving the following optimization problem:
$\min_{f,W,\{\xi_{e}\}_{e\in E}}\sum_{e\in E}\left(L(f(x)+W\xi_{e})+\lambda\|\xi_{e}\|^{2}\right),$ (19)
where $\lambda\|\xi_{e}\|^{2}$ is a regularization term. We used the implementation of CoDA provided on the authors’ GitHub site (https://github.com/yuan-yin/CoDA). The architecture is fixed, allowing only to change the various regularization coefficients as hyperparameters. For those we scanned similar values as the authors in their experiments, $\lambda_{\xi}\in\{1.e-2,1.e-3,1.e-4\}$, $\lambda_{\ell_{1}}\in\{1.e-5,1.e-6,1.e-7\}$ $\lambda_{\ell_{2}}\in\{1.e-5,1.e-6,1.e-7\}$, and $\text{dim}(\xi)\in\{3,6,10\}$. While overall working much better than LEADS, CoDA was still less reliable in recovering the long-term properties of the underlying systems than our method. Fig. 12 illustrates this for an example subject trained on the Lorenz-63, where the power spectrum learned by CoDA is much less accurate, especially in the low-frequency range.
@@ A.5 Unsupervised Feature Extraction
Table 5: Average classification accuracy of Gaussian Mixture Models (GMMs) applied to features extracted through various unsupervised methods on the EEG data. Values are across $10$ training runs for each model for non-deterministic approaches (hier-shPLRNN, MiniRocket and Attn-Conv-AE), with 10 randomly initialized GMMs evaluated per run.
@@ Appendix B Further Results
Figure 17: First PC of the 5d feature vector of a hier-shPLRNN, trained on the Lorenz-96 system in Sect. 4.1, aligns with the ground truth forcing parameter $F$ used to generate the training data.
Figure 18: Ground truth and freely generated arterial pressure wave (a), arterial flow velocity wave (b), arterial luminal area wave (c) and photoplethysmogram wave (d), simulated at $4$ different anatomical sites,for $10$ representative subjects (selected by k-medoids, color-coded).

### v2

@@ A.1.11 Hyperparameters
In Tab. 5 we report the hyperparameters used for training the shPLRNN models on the datasets in Sect. 4.1. These include the latent dimension $M$ (taken to be equivalent to the dimension of the observed data here, thus fixed), the hidden dimension $L$, the GTF interpolation factor $\alpha$, and the number of features $N_{\text{feat}}$. As recommended in Hess et al. [2023], $\alpha$ decays exponentially from $\alpha_{\text{start}}$ at the beginning to $\alpha_{\text{end}}$ at the end of training, where our values are based on Fig. S3 in that paper and not further fine-tuned (nor adaptively determined). The only parameters determined by grid search were thus $N_{\text{feat}}$ and $L$.
@@
Figure 16: Augmenting the hierarchical shPLRNN for identifying drivers of non-autonomous dynamics: Making system-specific features time-dependent. (a) A multi-layer perceptron (MLP), mapping time and subject-specific feature vectors $\bm{l}^{(j)}$ onto a time-dependent feature vector $\tilde{\bm{l}}^{(j)}_{t}$, is jointly trained with the hier-shPLRNN, resulting in explicitly time-dependent parameters of the DSR model. (b) To test this idea, we simulated a simple setup where the control parameter $\rho$ of the Lorenz-63 system was linearly increased in the range $[28,68]$, mimicking a time-dependent driving of the underlying system. For purposes of illustration, a hier-shPLRNN of just one subject was trained on time series segments from this simulation, but the approach could easily be extended to multiple subjects. (c) The temporal drift in $\rho$ was successfully recovered by the hier-shPLRNN, with a linear relation between actual and predicted $\rho$ ($R^{2}=0.996$). This simple example provides a proof of concept, but requires more detailed study to ensure this is a viable approach for more complex underlying systems and time dependencies in control parameters.
@@ A.2 Evaluation Measures
In lower dimensional state spaces this may be approximated by discrete binning: Let $m$ be the number of bins in each dimension and $\hat{p}^{(k)}$ the number of points in the $k$-th bin, then with a total number of bins $K=m^{d}$, an estimate of the state space divergence is given by
$D_{\text{stsp}}\simeq\sum_{k=1}^{K}\hat{p}_{\text{gen}}^{(k)}(x)\log\frac{\hat{p}_{\text{gen}}^{(k)}(x)}{\hat{p}_{\text{true}}^{(k)}(x)}$ (12)
Since the total number of bins scales exponentially with the observation dimension $N$, for experiments with the Lorenz-96 system, we used $m=5$ per dimension to ensure the approach remained computationally feasible. For very high-dimensional spaces, Brenner et al. [2022], Mikhaeil et al. [2022] suggest a GMM-based estimate with Gaussians centered on the trajectory points (which was not required here, however). For the values in Table 1, we sampled single time series with $10,000$ time steps from both the ground truth system (serving as the test set) and from each trained model, inferring only the initial state from the first time step of the test set, and cutting off transients ($1,000$ time steps) to focus on the long term attractor behavior.
@@ A.3 Datasets
Double Scroll (Chua’s Circuit)
The double-scroll attractor, often studied with Chua’s Circuit [Chua et al., 1986], is a 3-dimensional chaotic system. Its governing equations are given by:
$\dot{x}$ $=\alpha\left(y-x-f(x)\right),$ (18)
$\dot{y}$ $=x-y+z,$
$\dot{z}$ $=-\beta y,$
where the piecewise-linear function $f(x)$ is defined as:
$f(x)=m_{1}x+\frac{1}{2}(m_{0}-m_{1})\left(|x+1|-|x-1|\right).$ (19)
Standard parameter settings for a chaotic regime are $\alpha=9.0$, $\beta=14.0$, $m_{0}=-1.143$, and $m_{1}=-0.714$. For the results in Fig. 28, we chose $\alpha\in[8.4,9.4]$, leading to both single-scroll and double-scroll behavior.
@@
Human fMRI Data
The fMRI dataset analyzed in this study was initially collected by Koppe et al. [2014], re-analyzed in Koppe et al. [2019], and made publicly accessible as part of the latter work. It comprises time series from 10 brain regions bilaterally, chosen for their consistent involvement in n-back working memory tasks as identified in a meta-analysis by Owen et al. [2005]. Eigenvariates for these regions were extracted using the Brodmann area (BA) atlas. Specifically, the regions included medial and inferior parietal areas (BA7, BA40), prefrontal regions such as the dorsolateral prefrontal cortex (BA9, BA46) and ventrolateral prefrontal cortex (BA45, BA47), as well as the frontal pole (BA10), supplementary motor area (BA6), dorsal anterior cingulate cortex (BA32), and the cerebellum.
Electricity Load Data
The dataset from Trindade [2015] consists of electricity load time series from 370 sites recorded between 2011 and 2014. Each time series was preprocessed by removing leading zeros to account for differing start times across sites, resulting in varying maximum time lengths ($T_{\text{max}^{(j)}}$). The time series were standardized and slightly smoothed using a Gaussian kernel (window size = 5) and delay-embedded with a dimension of $d=5$ and a lag of $10$. We trained a feature vector with $N_{\text{feat}}=12$ for every site.
@@ A.4 Comparison Methods
As in LEADS, in CoDA (Context-Informed Dynamics Adaptation, Kirchmeyer et al. [2022]) the dynamics for each environment $e\in E$ is modeled as a combination of shared dynamics $f$ and environment-specific dynamics. Specifically, a low-dimensional context vector $\bm{\xi}_{e}$ is used to generate parameter variations via a hypernetwork, resulting in environment-specific parameters:
$\bm{\theta}_{e}=\bm{\theta}_{c}+\bm{W}\bm{\xi}_{e}$ (22)
where $\bm{\theta}_{c}$ are the shared parameters, $\bm{W}$ is a learned weight matrix, and $\bm{\xi}_{e}$ is the environment-specific context vector. The dynamics for each environment $e$ are then given by:
$f_{e}(x)=f(x;\bm{\theta}_{e})=f(x;\bm{\theta}_{c}+\bm{W}\bm{\xi}_{e}).$ (23)
As for LEADS, the learning objective is to minimize the contribution of the context vectors $\bm{\xi}_{e}$ while ensuring accurate modeling of the individual system dynamics. This is achieved by solving the following optimization problem:
$\min_{f,\bm{W},\{\bm{\xi}_{e}\}_{e\in E}}\sum_{e\in E}\left(L(f(x)+\bm{W}\bm{\xi}_{e})+\lambda\|\bm{\xi}_{e}\|^{2}\right),$ (24)
where $\lambda\|\bm{\xi}_{e}\|^{2}$ is a regularization term. We used the implementation of CoDA provided on the authors’ GitHub site (https://github.com/yuan-yin/CoDA). The architecture is fixed, allowing only to change the various regularization coefficients as hyperparameters. For those we scanned similar values as the authors in their experiments, $\lambda_{\bm{\xi}}\in\{1.e-2,1.e-3,1.e-4\}$, $\lambda_{\ell_{1}}\in\{1.e-5,1.e-6,1.e-7\}$ $\lambda_{\ell_{2}}\in\{1.e-5,1.e-6,1.e-7\}$, and $\text{dim}(\bm{\xi})\in\{3,6,10\}$. While overall working much better than LEADS, CoDA was still less reliable in recovering the long-term properties of the underlying systems than our method. Fig. 17 illustrates this for an example subject trained on the Lorenz-63, where the power spectrum learned by CoDA is much less accurate, especially in the low-frequency range.
@@ A.5 Unsupervised Feature Extraction
Table 7: Average classification accuracy of Gaussian Mixture Models (GMMs) applied to features extracted through various unsupervised methods on the EEG data. Values are across $10$ training runs for each model. A randomly initialized GMM was fitted for each training run, so variation in deterministic models stems from the different GMM fittings, while in ‘non-deterministic models’ (hier-shPLRNN, MiniRocket, Attn-Conv-AE) it comes from both training algorithm and GMM fitting.
@@
A.5.1 Hierarchical Markov Neural Operator
To illustrate the wider applicability of our approach, we also tested the hierarchization framework with the Markov Neural Operator (MNO) suggested in Li et al. [2022]. Code is publically available at https://github.com/neuraloperator/markov_neural_operator. As described in Sect. 3, we employed low-dimensional feature vectors and higher-dimensional group-level matrices linearly mapping onto all the trainable parameters of the MNO. We trained with $N_{feat}=1$ on the Lorenz-63 setup from sect. 4.2, and batch size, learning rates and initialization chosen as discussed in Appx. A.1. Following Li et al. [2022], we trained on one-step-ahead predictions, combined with a physics-informed dissipative loss. This loss enforces dissipativity by penalizing deviations from the global attractor through a regularization term that mimics energy dissipation. This initial implementation and the results in Fig. 18 are meant to highlight the straightforward applicability of the hierarchical framework to other architectures.
Figure 18: a: Ground truth and predicted $\rho$ from $1d$ feature values for the hierarchical MNO model (linear fit: $R^{2}=0.97$), similar to the results in Figs. 3 & 5. b: Long term freely generated dynamics for different $\rho$ values.
@@ Appendix B Further Results
Figure 23: First PC of the 5d feature vector of a hier-shPLRNN, trained on the Lorenz-96 system in Sect. 4.1 aligns with the ground truth forcing parameter $F$ used to generate the training data.
Figure 24: Ground truth (top) and freely generated (bottom) arterial pressure wave (a), arterial flow velocity wave (b), arterial luminal area wave (c) and photoplethysmogram wave (d), simulated at $4$ different anatomical sites, for $10$ representative subjects (selected by k-medoids, color-coded).
@@
Figure 27: a: Visualization of the first two PCs of the $12d$ inferred DS feature space for the electricity load data. Points represent sites, color-coded by cluster. Highlighted sites (black dots) correspond to examples shown below. Training data excerpts from three example sites illustrate pronounced differences in time series and highlight the highly nonstationary nature of the data. b: One-day test set forecast for each representative site, comparing ground truth (red) and model predictions (blue/orange/green). Predictions sometimes closely match ground truth values while capturing site-specific variations in complexity.
Figure 28: a: First 2 principal components (PCs) of the DS feature space for a model jointly trained on observations from the Lorenz-63 system, Rössler system, and Chua circuit, with $10$ different values each for the GT parameters $\rho$, $c$, and $\alpha$. The different DS are clearly separated along PC1, while PC2 distinguishes between the single-scroll versus double-scroll behavior of the Chua circuit (see c). b: Despite joint training on three different DS, the hier-shPLRNN embeds features in a low-dimensional space where the first four PCs capture almost all variation across subjects. c: Example reconstructions for all three systems.
Figure 29: Comparison of several popular time series (TS) models, N-HITS [Challu et al., 2022], Transformers [Vaswani et al., 2017], and Temporal Convolutional Networks (TCNs) [Bai et al., 2018], to the shPLRNN trained with GTF [Hess et al., 2023] with respect to forecasting performance (a) and long-term statistics (b, c) on observations ($T_{max}=5000$) from the chaotic Lorenz-63 attractor (Eq. 15). TS models were trained for 500 epochs using their implementations in the darts package [Herzen et al., 2022], with an input window size of 100 time steps, to forecast 30 time steps. All TS models had around $40,000$ parameters, while the shPLRNN only required $706$ trainable parameters in the chosen configuration ($M=3,L=100$). Long-term forecasts were generated – similar to the shPLRNN – by recursively feeding the TS models their own predictions. a: Short-term test set forecasts generated by the TSF models often match (or even exceed) the accuracy of the shPLRNN. The somewhat better forecasting performance is easily explained by the longer initialization windows used for the TS models compared to the shPLRNN (and potentially the much larger number of trainable parameters), which is initialized using just a single time step $\bm{x}_{0}$ at the beginning of the test set (in line with a state space’s Markov property in DS theory). b: However, unlike the shPLRNN, TS models fail to reconstruct the Lorenz-63 attractor geometry. c: Likewise, and unlike the shPLRNN, their long-term forecasts completely fail to capture the statistical temporal structure of the data, assessed here via the autocorrelation function [Wood, 2010] over the first $300$ time lags.